OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y =1/4 x + 2.

OpenStudy (anonymous):

y = 4x - 1 y = 4x - 3 y =1/4 x - 3 y = 1/4x - 1

OpenStudy (anonymous):

solve (y+2)=1/4(x+4)

OpenStudy (anonymous):

THANKS @begzat is that the answer?

OpenStudy (anonymous):

no you need to solve

OpenStudy (anonymous):

can you solve ?

OpenStudy (anonymous):

no its just im super confused in everything @begzat

OpenStudy (anonymous):

THANK YOU!!!